# Logical coherence in 2D compass codes

### v1

@@ Abstract
2D compass codes are a family of quantum error-correcting codes that contain the Bacon-Shor codes, the X-Shor and Z-Shor codes, and the rotated surface codes. Previous numerical results suggest that the surface code has a constant accuracy and coherence threshold under uniform coherent rotation. However, having analytical proof supporting a constant threshold is still an open problem. It is analytically proven that the toric code can exponentially suppress logical coherence in the code distance $L$. However, the current analytical lower bound on the threshold for the rotation angle $\theta$ is $|\sin(\theta)|<1/L$, which linearly vanishes in $L$ instead of being constant. We show that this lower bound is achievable by the Z-Shor code which does not have a threshold under stochastic noise. Compass codes provide a promising direction to improve on the previous bounds. We analytically determine thresholds for two new compass code families with thresholds near the rotated surface code’s numerically established coherence threshold. Furthermore, using a Majorana mode-based simulator, we use random families of compass codes to smoothly interpolate between the Z-Shor codes and the X-Shor codes.
@@
Coherent uniform rotation.— We simulate logical quantum memory instead of computation and assume perfect state preparation and syndrome extraction. We focus on the performance of encodings of a single logical qubit into a $[\![n,1,d]\!]$ error-correcting code under the channel that consists of coherent noise, X-syndrome measurement, and recovery. This physical channel acts on the physical state $\rho$ as $\mathcal{E}(\rho)=\sum_{s}Z(h_{s})\Pi_{s}\mathcal{N}_{\theta}(\rho)\Pi_{s}Z(h_{s})$, where $\Pi_{s}$ is the projector to the subspace compatible with measuring syndrome $s$, the recovery operator $Z(h_{s})$ is defined by applying Pauli Z operators on the support of the bit-string $h_{s}\in\mathbb{F}_{2}^{n}$, and $\mathcal{N}_{\theta}(\rho)$ is the uniform Z rotation by angle $\theta$ on all $n$ physical qubits, $\mathcal{N}_{\theta}(\rho)=U_{Z}\rho U_{Z}^{\dagger},\text{ with }U_{Z}=(e^{-i\frac{\theta}{2}Z})^{\otimes n}$.
@@
We can interpret the logical channel as a probabilistic mixture of logical rotations for $[\![n,1,d]\!]$ CSS codes with $k=1$ and $d$ odd [11, 7]. Following [7], for our Z-rotation-only case, we parameterize the Pauli Transfer Matrix (PTM) of the logical channel:
@@
where $\epsilon=\sum_{s}p_{s}(1-\cos(\theta_{s}))$ and $\delta=\sum_{s}p_{s}\sin(\theta_{s})$ are the weighted averages over $1-\cos$ and $\sin$ of the logical rotation angle $\theta_{s}$ for each syndrome $s$ and corresponding probability $p_{s}$. Intuitively, $\epsilon$ is a measure of average fidelity, and $\delta$ is a measure of coherence. The fact that we can define a logical angle $\theta_{s}$ for a given syndrome $s$ comes from the Pauli-expansion of $U_{Z}$, under the assumptions that there is only a single logical $\bar{Z}$ operator ($k=1$) and that all the codes in our discussion have even weight stabilizers and an odd distance, giving rise to an odd-weight logical $\bar{Z}$. For a derivation, see Lemma 4 in [11].
@@
Logical channel of the repetition code.— Consider a chain of qubits of length $d_{z}$. We define repetition codes (or phase flip codes) by their stabilizer group as $\mathcal{S}=\langle X_{i}X_{i+1}\mid i\in[1,d_{z}-1]\rangle$. Iverson and Preskill [7] calculated closed-form approximations for the logical channel parameters $\epsilon$ and $\delta$ as functions of $\theta$:
@@
where the approximation is based on Stirling’s formula, and valid when $\sin(\frac{\theta}{2})^{2}<1/2$, thus, when $|\theta|<\frac{\pi}{2}$. We heavily rely on these formulas below to understand quantum codes that can be decoded as products of repetition codes. As MWPM decoding is equivalent to ML decoding for repetition codes, our thresholds are considered to be under ML decoding.
2D compass codes.— 2D compass codes [32, 33] are CSS codes [34] closely related to the quantum compass model [35, 36], that describes interacting spins on an $l\times l$ lattice described by a Hamiltonian
@@
where a Pauli operator $P_{r,c}$ acts on qubits in row $r$ and column $c$. Compass codes are derived from the Bacon-Shor subsystem code [32, 37] based on $l^{2}$ physical qubits with a stabilizer group $S=\langle\prod_{r}X_{r,c}X_{r,c+1},\prod_{c}Z_{r,c}Z_{r+1,c}\rangle$ of $22l-$ 2 stabilizer generators and one logical qubit, leaving $l^{2}-2l+1$ gauge degrees of freedom in the gauge group $\mathcal{G}=\langle X_{r,c}X_{r,c+1},Z_{r,c}Z_{r+1,c}\rangle$, which are generated by the two-body terms of the compass model. Gauge operators $g\in\mathcal{G}$ can be “fixed” [38, 39] by moving $g$ to $S$ and removing all gauge operators from $\mathcal{G}$ that anti-commute with $g$. Compass codes are constructed through gauge fixing the Bacon-Shor code. In this paper, we also allow for rectangular compass codes; furthermore, we fix all gauge degrees of freedom, and thus our codes have $n-1$ stabilizer generators. We define compass codes on a $d_{x}\times d_{z}$ grid of qubits, $d_{z}$ odd (even distance codes have no coherence in their logical channel). Three known families of codes are the Z-Shor codes/X-Shor codes and the rotated surface code [26]. It is useful to think of these codes as red (Z-type) and blue (X-type) colorings of a checkerboard [33].
Coherence threshold for the Z-Shor code.— We first introduce a new measure for coherence. To leading order, the growth of $r_{m}$ in $m$ is linear for stochastic channels but quadratic for purely coherent channels, as shown by Iverson and Preskill [7]. For the purely coherent channel, the approximate expression for $r_{m}$ (when $\frac{m^{2}\epsilon}{3}$ is small) shows that $\delta$ drives the quadratic term: $r_{m}=\frac{1}{3}\epsilon m-\frac{1}{6}m(m-1)(\epsilon^{2}-\delta^{2})+\mathcal{O}(m^{3}\epsilon^{3},m^{3}\epsilon\delta^{2},m^{4}\delta^{4})$. When $\epsilon$ is small, the $\epsilon^{2}$ term can be ignored, and it is justifiable to choose $\kappa(\theta)\equiv\frac{\delta(\theta)^{2}}{\epsilon(\theta)}$ as a simple measure for coherence representing the strength of the quadratic term in $r_{m}$ relative to the linear term. Our measure of coherence is then for the repetition codes based on the approximate $\epsilon$ and $\delta$ values by Iverson and Preskill:
@@
For any $\theta<\frac{\pi}{2}$, both $\kappa_{l}(\theta)$ and $\epsilon_{l}(\theta)$ decrease exponentially in $l$, thus we conclude that the coherence threshold and the fidelity threshold are both $\theta_{th}^{rep}=\frac{\pi}{2}$ for the repetition code.
Coherence threshold for the Z-Shor code.— The $d_{x}=1$ Z-Shor code is the length $d_{z}$ phase-flip code. For $d_{x}>1$, the fact that any logical state $|\psi_{L}\rangle$ is stabilized $Z_{i,j}Z_{i+1,j}$ causes the property that $e^{-i\frac{\theta}{2}Z_{i,j}}|\psi_{L}\rangle=e^{-i\frac{\theta}{2}Z_{i+1,j}}|\psi_{L}\rangle$. Using this, in every column $j$, every single qubit Z-rotation on row $i$ can be equivalent to acting on qubit $1,j$, resulting in an effective rotation of $e^{-id_{x}\frac{\theta}{2}Z_{1,j}}$, which we refer to as “zipping” of the rotation along ZZ stabilizers.
@@
Coherence threshold for the X-Shor code.— We can think of the X-Shor code as $d_{x}$ length-$d_{z}$ phase-flip codes glued together by weight-$2d_{z}$ Z operators. As such, the logical channel of the X-Shor code is equivalent to applying $d_{x}$ times the logical channel of the length-$d_{z}$ phase-flip code. In the Pauli Transfer Matrix representation, the combined PTM of applying multiple channels is simply the product of the individual channel PTMs. In our case, the PTM for the X-Shor code’s logical channel is going to be the product of the PTMs of the phase-flip code logical channel.
@@
where the approximations are based on first-order expansions in $\epsilon_{d_{z}},\lambda_{d_{z}}$ around zero, and thus they apply when these values are sufficiently small. In this low-noise regime, by further ignoring negligible third and second-order terms, we find that $\kappa_{X,d_{x}\times d_{z}}(\theta)\approx d_{x}\kappa_{d_{z}}(\theta)$, which means that the X-Shor code has the same coherence threshold as the phase-flip code. Thus, for $\theta$ values below $\theta_{th}=\pi/2$, the X-Shor code suppresses loss of fidelity and coherence of the logical channel exponentially in $d_{z}$. This suppression is tempered by a factor of $d_{x}$, which, in the case of square codes $l=d_{x}=d_{z}$, is a linear slowdown.
FIGURE 1: Two members of the Z-stacked Shor code family, $C_{7,2}$ (a) and $C_{7,3}$ (b). X stabilizers are blue rectangles between two columns of qubits spanning from top to bottom, while Z stabilizers are red rectangles between two rows of qubits spanning from left to right. In (b), the second Z-Shor block shows the uniform Z rotation by $\theta$. The top block shows the “zipped” view, which is equivalent to the uniform rotation up to the weight-2 Z-stabilizers. Here, qubits in the first row experience a $3\theta$ rotation while the other two rows do not participate.
Coherence threshold for the Z-stacked Shor codes.— We now define a generalization of the X-Shor codes, the family of Z-stacked Shor codes, with members of length $l$ and block-height $h$, denoted $\mathcal{C}_{l,h}$, with $\mathcal{C}_{l,1}$ being the family of $l\times l$ square X-Shor codes. Consider the $l\times l$ square grid of qubits. In the first $l_{h}\equiv\lfloor l/h\rfloor$ rows, we glue together $h\times l$ Z-Shor code blocks with weight-$2l$ $Z$ stabilizer generators. In the remaining $t_{h}\equiv l\mod h$ rows, we glue together $1\times l$ Z-Shor code blocks (which are phase flip codes) with the $Z$ stabilizer generators. Two $l=7$ examples for $h=2$ and $h=3$ are displayed in Fig. 1. As the structure of these codes is similar to the X-Shor code, the logical channel’s PTM is the product of the Z-Shor code PTMs. From before, we know that for each Z-Shor code block, we can use the parameters of the phase-flip code, just replacing $\theta$ with $h\theta$. Thus, the logical channel parameters for $\mathcal{C}_{l,h}$ are:

### v2

@@ Abstract
2D compass codes are a family of quantum error-correcting codes that contain the Bacon-Shor codes, the $X$-Shor and $Z$-Shor codes, and the rotated surface codes. Previous numerical results suggest that the surface code has a constant accuracy and coherence threshold under uniform coherent rotation. However, having analytical proof supporting a constant threshold is still an open problem. It is analytically proven that the toric code can exponentially suppress logical coherence in the code distance $L$. However, the current analytical lower bound on the threshold for the rotation angle $\theta$ is $|\sin(\theta)|<1/L$, which linearly vanishes in $L$ instead of being constant. We show that this lower bound is achievable by the $Z$-Shor code which does not have a threshold under stochastic noise. Compass codes provide a promising direction to improve on the previous bounds. We analytically determine thresholds for two new compass code families that provide upper and lower bounds to the rotated surface code’s numerically established infidelity threshold. Furthermore, using a Majorana mode-based simulator, we use random families of compass codes to smoothly interpolate between the $Z$-Shor codes and the $X$-Shor codes.
I Introduction
@@
The paper is structured as follows. In Section II, we establish notation and review the logical channel under uniform coherent rotation and review 2D compass codes. In Section III we derive coherence thresholds for repetition code like compass code families and showcase our numerical results. Finally, we conclude our findings in Section IV.
II Background
We simulate logical quantum memory instead of computation and assume perfect state preparation and syndrome extraction. We focus on the performance of encodings of a single logical qubit into a $[\![n,1,d]\!]$ error-correcting code under the channel that consists of coherent noise, $X$-syndrome measurement, and recovery. This physical channel acts on the physical state $\rho$ as
$\mathcal{E}(\rho)=\sum_{s}Z(h_{s})\Pi_{s}\mathcal{N}_{\theta}(\rho)\Pi_{s}Z(h_{s}),$ (1)
where $\Pi_{s}$ is the projector to the subspace compatible with measuring syndrome $s$, the recovery operator $Z(h_{s})$ is defined by applying Pauli Z operators on the support of the bit-string $h_{s}\in\mathbb{F}_{2}^{n}$, and $\mathcal{N}_{\theta}(\rho)$ is the uniform Z rotation by angle $\theta$ on all $n$ physical qubits, $\mathcal{N}_{\theta}(\rho)=U_{Z}\rho U_{Z}^{\dagger},\text{ with }U_{Z}=(e^{-i\frac{\theta}{2}Z})^{\otimes n}$.
We are interested in how much the logical channel differs from the identity channel and how this difference grows by applying the logical channel repeatedly $m$ times. It has been previously shown that in the limit of large code sizes, the logical channel of stabilizer codes is close to a stochastic channel, thus well characterized by the logical infidelity [6], and is worthwhile optimizing decoding for. Note that this is in contrast with physical channels, where gate fidelity might not be predictive of scalability [28, 29]. Our distance metric is the average infidelity of the logical states after $m$ applications of the logical channel, which we simply refer to as the infidelity of the logical channel:
@@ II Background
We can interpret a single application of the logical channel as a probabilistic mixture of logical rotations for $[\![n,1,d]\!]$ CSS codes with $k=1$ and $d$ odd [11, 7]. Following [7], for our $Z$-rotation-only case, we parameterize the Pauli Transfer Matrix (PTM) of the logical channel:
@@
where $\epsilon=\sum_{s}p_{s}(1-\cos(\theta_{s}))$ and $\delta=\sum_{s}p_{s}\sin(\theta_{s})$ are the averages of $1-\cos(\theta_{s})$ and $\sin(\theta_{s})$ over the distribution of logical rotation angles $\theta_{s}$ for syndromes $s$ with corresponding probabilities $p_{s}$. For a Hilbert-space with dimension $d=2^{n}$ for $n$ qubits, the PTM channel representation follows from representing the density matrix as a vector $\rho=\sum_{j=0}^{d^{2}-1}\rho_{j}\sigma^{j}$ in the Pauli basis $\{\rho^{j}\}$. Then the channel becomes a matrix $\bar{N}_{ij}=\text{Tr}[\mathcal{N}(\sigma^{j})\sigma^{i}]$.
The fact that we can define a logical angle $\theta_{s}$ for a given syndrome $s$ comes from the Pauli-expansion of $U_{Z}$, under the assumptions that there is only a single logical $\bar{Z}$ operator ($k=1$) and that all the codes in our discussion have even weight stabilizers and an odd distance, giving rise to an odd-weight logical $\bar{Z}$. For a derivation, see Lemma 4 in [11].
For a single application of the logical channel, $r_{1}=\frac{\epsilon}{3}$ [7], thus $\epsilon$ is proportional to infidelity. A non-zero $\delta$ indicates rotation between logical $X$ and $Y$. For example, when a Pauli channel is completely incoherent, then its PTM is diagonal, with $\delta=0$. For a completely coherent, single qubit rotation $U=e^{-i\theta/2Z}$, the diagonal element is $1-\epsilon=\cos(\theta)$, and the off-diagonal element is $\delta=\sin(\theta)$. As we’ll see in Section III, only $\epsilon$ contributes to $r_{1}$, but both of these terms contribute to $r_{m}$. To leading order, $\epsilon$ contributes linearly, while $\delta$ contributes a quadratic term in $m$.
@@
II.1 Logical channel of the repetition code
Consider a chain of qubits of length $d_{z}$. We define repetition codes (or phase flip codes) by their stabilizer group as $\mathcal{S}=\langle X_{i}X_{i+1}\mid i\in[1,d_{Z}-1]\rangle$. Iverson and Preskill [7] calculated closed-form approximations for the logical channel parameters $\epsilon$ and $\delta$ as functions of $\theta$:
@@ II.1 Logical channel of the repetition code
where the approximation is based on Stirling’s formula, and valid when $\sin(\frac{\theta}{2})^{2}<1/2$, thus, when $|\theta|<\frac{\pi}{2}$. We heavily rely on these formulas below to understand quantum codes that can be decoded as products of repetition codes.
II.2 2D compass codes
2D compass codes [32, 33] are CSS codes [34] closely related to the quantum compass model [35, 36], that describes interacting spins on an $l\times l$ lattice described by a Hamiltonian
@@
where a Pauli operator $P_{r,c}$ acts on qubits in row $r$ and column $c$. Compass codes are derived from the Bacon-Shor subsystem code [32, 37] based on $l^{2}$ physical qubits with a stabilizer group $S=\langle\prod_{r}X_{r,c}X_{r,c+1},\prod_{c}Z_{r,c}Z_{r+1,c}\rangle$ of $22l-$ 2 stabilizer generators and one logical qubit, leaving $l^{2}-2l+1$ gauge degrees of freedom in the gauge group $\mathcal{G}=\langle X_{r,c}X_{r,c+1},Z_{r,c}Z_{r+1,c}\rangle$, which are generated by the two-body terms of the compass model. Gauge operators $g\in\mathcal{G}$ can be “fixed” [38, 39] by moving $g$ to $S$ and removing all gauge operators from $\mathcal{G}$ that anti-commute with $g$. Compass codes are constructed through gauge fixing the Bacon-Shor code. In this paper, we also allow for rectangular compass codes; furthermore, we fix all gauge degrees of freedom, and thus our codes have $n-1$ stabilizer generators. We define compass codes on a $d_{x}\times d_{z}$ grid of qubits, $d_{z}$ odd (even distance codes have no coherence in their logical channel). Three known families of codes are the $Z$-Shor codes, $X$-Shor codes, and the rotated surface code [26]. We call Shor’s code [3] the $Z$-Shor code, to emphasize the fact that the weight-2 stabilizer generators are $Z$-type, displayed in Fig. 1(a). The Hadamard transform of the $Z$-Shor code is the $X$-Shor code, where the weight-2 stabilizers are $X$-type, displayed in Fig. 1(b). It is useful to think of 2D compass codes as red ($Z$-type) and blue ($X$-type) colorings of a $(d_{X}-1)\times(d_{Z}-1)$ size grid [33], in which case, $Z$-Shor is all blue, and $X$-Shor is all red and the rotated surface code follows a checkerboard pattern.
(a) $Z$-Shor code
(b) $X$-Shor code
FIGURE 1: (color online) (a) The 5x5 Shor’s code which we call $Z$-Shor code, referring to the type of the weight-2 stabilizer generators and (b) the 5x5 $X$-Shor code.
III Analytical and numerical coherence thresholds
III.1 Coherence threshold for the repetition code
We first introduce a new measure for coherence. To leading order, the growth of $r_{m}$ in $m$ is linear for stochastic channels but quadratic for purely coherent channels, as shown by Iverson and Preskill [7]. For the repetition code’s logical channel, Iverson and Preskill derived the approximate expression for $r_{m}$ (when $\frac{m^{2}\epsilon}{3}$ is small), which shows that $\delta$ drives the quadratic term:
$r_{m}$ $=\frac{1}{3}\epsilon m-\frac{1}{6}m(m-1)\delta^{2}$ (5)
$+\mathcal{O}(m^{2}\epsilon^{2},m^{3}\epsilon^{3},m^{3}\epsilon\delta^{2},m^{4}\delta^{4})$ (6)
It is now clear why $\epsilon$ is a measure of infidelity, as for a single repetition, $r_{1}\approx\frac{1}{3}\epsilon$. Thus, we choose $\kappa(\theta)\equiv\frac{\delta(\theta)^{2}}{\epsilon(\theta)}$ as a simple measure for coherence representing the strength of the quadratic term in $r_{m}$ relative to the linear term. Then, the coherence of the logical channel of repetition codes, after plugging in the approximate $\epsilon$ and $\delta$ values described by Eqs. 3 and 4 is as follows:
@@ III.1 Coherence threshold for the repetition code
We define the coherence threshold $\theta_{th}$ to be the rotation value under which it is guaranteed that $\kappa_{l}(\theta)$ is suppressed exponentially in the code size $l$. The infidelity threshold $\theta_{th}^{r_{1}}$ is the rotation value for which all $\theta\leq\theta_{th}^{r_{1}}$, $r_{1}(\theta)$ is exponentially suppressed in the code size $l$. For any $\theta<\frac{\pi}{2}$, both $\kappa_{l}(\theta)$ and $\epsilon_{l}(\theta)$ decrease exponentially in $l$, thus we conclude that the coherence threshold and the infidelity threshold are both $\theta_{th}^{rep}=\frac{\pi}{2}$ for the repetition code.
As minimum-weight decoding is equivalent to ML decoding for repetition codes under code capacity noise models, these repetition code thresholds can be considered to be both under ML and minimum-weight decoding. See Appendix B for further numerical evidence on this.
III.2 Coherence threshold for the $Z$-Shor code
The $d_{x}=1$ $Z$-Shor code is the length $d_{z}$ phase-flip code. For $d_{x}>1$, the fact that any logical state $|\psi_{L}\rangle$ is stabilized $Z_{i,j}Z_{i+1,j}$ causes the property that $e^{-i\frac{\theta}{2}Z_{i,j}}|\psi_{L}\rangle=e^{-i\frac{\theta}{2}Z_{i+1,j}}|\psi_{L}\rangle$. Using this, in every column $j$, every single qubit $Z$-rotation on row $i$ can be equivalent to acting on qubit $1,j$, resulting in an effective rotation of $e^{-id_{x}\frac{\theta}{2}Z_{1,j}}$, which we refer to as “zipping” of the rotation along ZZ stabilizers.
@@ III.2 Coherence threshold for the $Z$-Shor code
Physical rotations above $\theta=\frac{\pi}{2d_{x}}$ for the $Z$-Shor code are equivalent to being above the repetition code threshold of $\pi/2$, and thus the the minimum weight recovery becomes incorrect. An ML decoder [40] would then find the correct recovery operator. Thus, in the case of the $Z$-Shor code, ML and minimum weight recovery are not the same and our repetition code-based threshold for the $Z$-Shor code should be interpreted to be under minimum-weight recovery. In Appendix B we provide further numerical evidence for this argument.
III.3 Coherence threshold for the $X$-Shor code
We can think of the $X$-Shor code as $d_{x}$ length-$d_{z}$ phase-flip codes glued together by weight-$2d_{z}$ Z operators. As such, the logical channel of the $X$-Shor code is equivalent to applying $d_{x}$ times the logical channel of the length-$d_{z}$ phase-flip code. In the Pauli Transfer Matrix representation, the combined PTM of applying multiple channels is simply the product of the individual channel PTMs. In our case, the PTM for the $X$-Shor code’s logical channel is going to be the product of the PTMs of the phase-flip code logical channel.
@@ III.3 Coherence threshold for the $X$-Shor code
where the approximations are based on first-order expansions in $\epsilon_{d_{z}},\lambda_{d_{z}}$ around zero, and thus they apply when these values are sufficiently small. In this low-noise regime, by further ignoring negligible third and second-order terms, we find that $\kappa_{X,d_{x}\times d_{z}}(\theta)\approx d_{x}\kappa_{d_{z}}(\theta)$, which means that the $X$-Shor code has the same coherence threshold as the phase-flip code. Thus, for $\theta$ values below $\theta_{th}=\pi/2$, the $X$-Shor code suppresses infidelity and coherence of the logical channel exponentially in $d_{z}$. This suppression is tempered by a factor of $d_{x}$, which, in the case of square codes $l=d_{x}=d_{z}$, is a linear slowdown.
In the case of the $X$-Shor code, our calculations reflect using a combined mininum weight recovery for each row of repetition codes, thus the threshold above is under MWPM decoding (see Appendix B for numerical evidence).
FIGURE 2: (color online) Two members of the $Z$-stacked Shor code family, $C_{7,2}$ (a) and $C_{7,3}$ (b). X stabilizers are blue (lighter) rectangles between two columns of qubits spanning from top to bottom, while Z stabilizers are orange (darker) rectangles between two rows of qubits spanning from left to right. In (b), the second $Z$-Shor block shows the uniform Z rotation by $\theta$. The top block shows the “zipped” view, which is equivalent to the uniform rotation up to the weight-2 $Z$-stabilizers. Here, qubits in the first row experience a $3\theta$ rotation while the other two rows do not participate.
III.4 Coherence threshold for the $Z$-stacked Shor codes
We now define a generalization of the $X$-Shor codes, the family of $Z$-stacked Shor codes, with members of length $l$ and block-height $h$, denoted $\mathcal{C}_{l,h}$, with $\mathcal{C}_{l,1}$ being the family of $l\times l$ square $X$-Shor codes. Consider the $l\times l$ square grid of qubits. In the first $l_{h}\equiv\lfloor l/h\rfloor$ rows, we glue together $h\times l$ $Z$-Shor code blocks with weight-$2l$ $Z$ stabilizer generators. In the remaining $t_{h}\equiv l\mod h$ rows, we glue together $1\times l$ $Z$-Shor code blocks (which are phase flip codes) with the $Z$ stabilizer generators. Two $l=7$ examples for $h=2$ and $h=3$ are displayed in Fig. 2. As the structure of these codes is similar to the $X$-Shor code, the logical channel’s PTM is the product of the $Z$-Shor code PTMs. From before, we know that for each $Z$-Shor code block, we can use the parameters of the phase-flip code, just replacing $\theta$ with $h\theta$. Thus, the logical channel parameters for $\mathcal{C}_{l,h}$ are: